# A Calculus of Overlays

### v1

@@ 4. Translation from $\lambda$-Calculus
A $\lambda$-bound variable $x$ with de Bruijn index $n$ (the number of enclosing $\lambda$-abstractions between the reference and its binder) becomes $\uparrow^{\!n}.\mathrm{argument}$. The indexed reference $\uparrow^{\!n}$ traverses $n$ scope levels upward to reach the abstraction that binds $x$, giving access to the $\mathrm{argument}$ slot after all inheritance. A let-bound variable $x$ becomes the lexical reference $x.\mathrm{result}$, projecting the application result from the sibling property $x$ in the enclosing record.
@@
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to Overlay-Calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is Overlay-Calculus; Appendix A further shows that Overlay-Calculus primitives directly encode a random-access machine.
@@
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices—no function spaces, no continuity conditions, no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix E), which requires only set-theoretic monotonicity—strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree—no substitution occurs at any point.
@@
That a semantics of the $\lambda$-calculus can be given without function spaces is perhaps unexpected: function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function. Section 4.2 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate (a term converges iff it has a head normal form) and fully abstract (two terms are overlay-contextually equivalent iff they have the same Böhm tree).
Combining these two facts—elementary machinery and full abstraction—yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Hyland and Ong’s game semantics (Hyland and Ong, 2000) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
The reverse direction is non-trivial, as we discuss next.
4.1. Expressive Asymmetry
The translation above consists of five rules; the full computational power of the $\lambda$-calculus embeds into Overlay-Calculus with no additional machinery.
The reverse direction—encoding Overlay-Calculus in pure $\lambda$-calculus—is non-trivial. Defining the meaning of Overlay-Calculus requires machinery for lazy allocation, recursive data structures, and set-valued resolution that have no direct counterparts in the $\lambda$-calculus. This sophistication is not accidental—it reflects three features:
  1. (1)
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
  3. (3)
    Commutative, idempotent merge. Inheritance in Overlay-Calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery.
Overlay-Calculus is not the $\lambda$-calculus in disguise: it provides primitive operations for computational patterns that are derivable but complex in the $\lambda$-calculus.
The reverse asymmetry also exists: nested expressions such as f(gx)f\;(g\;x) are primitive in the $\lambda$-calculus but require manual A-normal form conversion in Overlay-Calculus—the intermediate result gxg\;x must be given an explicit name. However, ANF conversion is a trivial, mechanical transformation (Flanagan et al., 1993); a surface language built on Overlay-Calculus can support nested expressions as syntactic sugar without extending the calculus itself.
@@ 4.2. Böhm Tree Correspondence
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties—the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
@@
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984)—the canonical fully abstract model of the lazy $\lambda$-calculus—using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix F.
5. Case Study: The Expression Problem
The Expression Problem (Wadler, 1998) asks how to extend a data type with both new cases and new operations, without modifying existing code. Existing solutions require object algebras (d. S. Oliveira and Cook, 2012; Wang and d. S. Oliveira, 2016), finally tagless interpreters (Carette et al., 2009), or type-class machinery. In a typed setting the problem additionally requires static type safety; Overlay-Calculus is untyped, so we focus on the extensibility aspect. Suppose a standard library provides a $\mathrm{binnat}$ module for binary natural arithmetic (see Appendix B). We define an expression language with an evaluation operation:
$\{\quad\mathrm{expression}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{magnitude}\mapsto\{\}\},\;\mathrm{Addition}\mapsto\{\mathrm{left}\mapsto\{\},\;\mathrm{right}\mapsto\{\}\}\},$
$\mathrm{evaluation}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{outcome}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\qquad\mathrm{Addition}\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
@@ 5. Case Study: The Expression Problem
$\qquad\qquad\mathrm{left\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{outcome},$
@@
$\qquad\qquad\mathrm{right\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{outcome},$
$\qquad\qquad\mathrm{sum}\mapsto\{\mathrm{binnat}.\mathrm{Addition},\;\mathrm{augend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{left\_outcome},\;\mathrm{addend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{right\_outcome}\},$
$\qquad\qquad\mathrm{outcome}\mapsto\mathrm{sum}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Addition}\mapsto\{\mathrm{augend}\mapsto\{\},\;$
$\qquad\qquad\mathrm{addend}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}\quad\}$
A new operation (e.g., display) is added by defining a new module that inherits from the same $\mathrm{expression}$ schemas:
$\mathrm{display}\mapsto\{\mathrm{Constant}$ $\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{representation}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\mathrm{Addition}$ $\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
$\qquad\mathrm{representation}\mapsto\{$
$\qquad\qquad\mathrm{left}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{representation},\;$
$\qquad\qquad\mathrm{right}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{representation}\}\}\}$
A new case (e.g., negation) is added with its operation handlers and an empty API for the arithmetic primitive:
$\mathrm{negation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\}\}\},$
$\mathrm{negation\_evaluation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{negation}.\mathrm{Negation},\;$
$\qquad\mathrm{operand\_outcome}\mapsto\mathrm{Negation}.\mathbf{this}.\mathrm{operand}.\mathrm{outcome},$
$\qquad\mathrm{negated}\mapsto\{\mathrm{binnat}.\mathrm{Negation},\;\mathrm{operand}\mapsto\mathrm{operand\_outcome}\},$
$\qquad\mathrm{outcome}\mapsto\mathrm{negated}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}$
All modules compose freely; none requires modification of existing definitions. A complete executable version of this example—including both the evaluation and display operations, a negation extension, and their free inheritance—is available in the implementation (Yang, 2025b), along with test cases verifying evaluation and display for nested expressions mixing old and new cases.
Note that $\mathrm{magnitude}\mapsto\{\}$ in $\mathrm{Constant}$ resembles a type annotation but is purely structural in Overlay-Calculus: the evaluator works as long as the required properties appear somewhere in the inheritance chain, regardless of any schema declarations.

### v2

@@ 4. Embedding the $\lambda$-Calculus
[…]
@@
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to overlay-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is overlay-calculus.
@@
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
@@
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract, meaning two terms are overlay-contextually equivalent iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
@@ 4.1. Böhm Tree Correspondence
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
@@
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984), the canonical fully abstract model of the lazy $\lambda$-calculus, using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in overlay-calculus by overlaying new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), overlay-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: overlay definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as overlay files (Yang, 2025b). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
NatData
Following the Church-encoding convention (Yang, 2025b), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
$\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Product}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\}\},$
$\mathrm{Nat}\mapsto\{\mathrm{NatFactory},\;\mathrm{NatFactory}.\mathrm{Product}\}\}$
NatPlus
Addition is a separate overlay file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
$\mathrm{NatPlus}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{sum}\mapsto\{\mathrm{Product}\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},\;\mathrm{sum}\mapsto\{\mathrm{addend}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},$
$\qquad\mathrm{\_increasedAddend}\mapsto\{\mathrm{Successor},\;\mathrm{predecessor}\mapsto\{\mathrm{addend}\}\},$
@@ 5. Case Study: Nat Arithmetic
$\qquad\mathrm{\_recursiveAddition}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Plus},$
$\qquad\quad\mathrm{addend}\mapsto\{\mathrm{\_increasedAddend}\}\},$
$\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
NatEquality
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the constructor; we use the Visitor pattern. $\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
$\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
$\mathrm{BooleanData}$ defines the output sort:
$\mathrm{BooleanData}\mapsto\{\;$ $\mathrm{BooleanFactory}\mapsto\{\mathrm{Product}\mapsto\{\},\;\mathrm{True}\mapsto\{\mathrm{Product}\},\;\mathrm{False}\mapsto\{\mathrm{Product}\}\},$
$\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
$\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
$\mathrm{NatFactory}\mapsto\{$
$\qquad\mathrm{Product}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
$\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
$\hskip 18.49988pt\mathrm{RecursiveEquality}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Equal},$
$\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
$\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{One}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Zero}\}\},$
$\mathrm{Two}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{One}\}\},$
$\mathrm{Three}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Two}\}\},$
$\mathrm{Four}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Three}\}\},$
$\mathrm{Five}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Four}\}\}\}$
The arithmetic test inherits $\mathrm{NatConstants}$ together with the operations it needs:
$\mathrm{Test}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
@@
$\mathrm{Addition}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Two}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Three}\}\},$
$\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$. No existing file was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
Tree-level inheritance and return types
Each operation is an independent overlay that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
Church-encoded Nats are tries
A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
$\mathrm{CartesianTest}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{OneOrTwo}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{One},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Two}\},$
$\mathrm{ThreeOrFour}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Three},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Four}\},$
$\mathrm{Result}\mapsto\{\mathrm{OneOrTwo}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{ThreeOrFour}\}\},$
$\mathrm{Check}\mapsto\{\mathrm{Result}.\mathrm{sum}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Result}.\mathrm{sum}\}\}\}$
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overlays}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.